OpenStudy (anonymous):

3. You research the average cost of a movie ticket over several years to look for trends. The table below shows your data. What is the equation for a line of best fit? Approximately how much would you expect to pay for a movie ticket in 2021? Let x be the number of years after 1980. Year: 1980, 1984, 1988, 1992, 1996, 2000, 2004, 2008, 2012. Price: $4.00, $4.50, $5.25, $5.50, $7.00, $7.75, $8.25, $8.78, $9.00. Answer choices: A. y=0.17x+4; $11 B. y=4x; $164 C. y=5.7x+4; $238 D. y=0.67x-4; $23

OpenStudy (31356):

3) Given Year: 1980, 1984, 1988, 1992, 1996, 2000, 2004, 2008, 2012 Price: $4.00, $4.50, $5.25, $5.50, $7.00, $7.75, $8.25, $8.78, $9.00 Line of best fit Ticket in 2021 Let x be 1980 The formula for the linear regression is y = ax + b where a (slope) = [n(Σxy) - (Σx)(Σy)]/[n(Σx^2) - (Σx)^2] b (intercept) = [Σy - a(Σx)]/n n = 9 Σx = 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 = 45 Σy = 4.00 + 4.50 + 5.25 + 5.50 + 7.00 + 7.75 + 8.25 + 8.78 + 9.00 = 60.03 Σx^2 = 1^2 + 2^2 + 3^2 + 4^2 + 5^2 + 6^2 + 7^2 + 8^2 + 9^2 = 285 Σxy = 1*4 + 2*4.5 + 3*5.25 + 4*5.5 + 5*7 + 6*7.75 + 7*8.25 + 8*8.78 + 9*9 = 341.24 a = [9(341.24) - (45)(60.03)]/[9(285) - (45)^2] = 0.6848 b = [60.03 - 0.6848(45)]/9 = 3.246 The equation for a line of best fit is y = 0.6848x + 3.246 In 2021 (2021 - 1979 = 42) the ticket price would be y = 0.6848*42 + 3.246 = 32 Answer: It is close to D but not quite the same. Perhaps it is typographical?

OpenStudy (31356):

Medal?

OpenStudy (31356):

Thanks!

OpenStudy (anonymous):

np:) thx